View History Bookmarks Tools Window Help 44%E) wed ECON1000 Assignment I(CMU Wn X T7058ECON1000 Assignment l (CMU Winter 20 回 ☆ ofl9 - + Automatic Zoom 2. The next four questions refer to the following price and output data over a five-year period for an economy that produces only one good. Assume that year 2 is the base year Units ofPrice Year Output per unit 16 20 30 36 40 $2 (a) If year 2 is the base year, give the price index for year 3. (b) Give the nominal GDP for year 4. (c) What is the real GDP for year 4? (d) Tell which years you would deflate nominal GDP and which years you would inflate nominal GDP in finding real GDP Book Pro 名。F3 2 3 4
Answer #1

a)

Price in year 2=$3

Price in year 3=$4

Price index for year 3=(Price in year 3/Price in year 2)*100 =133.33

b)

Nominal GDP for Year 4 will be equal to value of goods purchased at year 4 price.

Nominal GDP=Price*Units of output=5*36=$180

c)

Price in year 2=$3

Price in year 4=$5

Price index for year 4=Price in year 4/Price in year 2*100=(5/3)*100=166.6667

Real GDP will be equal to the value of output at base year prices=(Nominal GDP/Price index for year 4)*100=(180/166.6667)*100=108

Now, we calculate the value of output at base year prices.

Real GDP for year 4=36*3=$108

We reached at same figures.

d)

Let us calculate Price Index for each year

Price index for year 1=(Price in year 1/Price in year 2)*100=(2/3)*100=66.67

Price index for year 2=(Price in year 2/Price in year 2)*100=(3/3)*100=100

Price index for year 3=(Price in year 3/Price in year 2)*100=(4/3)*100=133.33

Price index for year 4=(Price in year 4/Price in year 2)*100=(5/3)*100=166.67

Price index for year 5=(Price in year 5/Price in year 2)*100=(6/3)*100=200

For finding the real GDP, GDP will be inflated for years which have price index value less than 100 i.e year 1.

For finding the real GDP, GDP will be deflated for years which have price index value higher than 100 i.e. year 3,4 and 5.

Real GDP will be same as nominal GDP in case of base year.